The length of a rectangle is six times its width.
If the area of the rectangle is
600m^2 find its perimeter.

Answer by ewatrrr(24785)

 
Hi,
.
(6w)w = 600m^2
w^2 = 100m^2
w = 10m (Using positive solution for unit measure)
L = 60m
P = 2*10m +2*60m = 140m